If I deposit ten dollars in saving which paid two percent monthly, for fifty years, how much money would I have.
Answer by oscargut(2103)

Solution:
Amount = 10(1.02)^600 = 1,445,782.81 (I used compound interest)
Answer: 1,445,782.81
If you have any doubt or if you have more problems my e-mail is : mthman@gmail.com